1. A polarized light of intensity 10 is passed through another polarizer whose pass axis makes an angke of 60 degrees with the pass axis of the former. What is the intensity of emergent polarized light from second polarizer?

2. What angle is needed between the direction of polarized light and the axis of a polarizing filter to reduce its intensity by 90.0%?


1
Expert's answer
2021-06-07T09:34:26-0400

1. According to the Malus's law, the intensity of emergent polarized light from second polarizer is given as follows:


"I = I_0\cos\theta"

where "I_0 = 10" is the initial light intensity, "\theta = 60\degree" is the angle between the polarizers. Thus, obtain:


"I = 10\cdot \cos60\degree = 10\cdot 0.5 = 5"

2. If the intensity is rediced by 90.0%, than "I\/I_0 = 0.9". Expressing the angle from the Malus's law, obtain:


"\theta = \arccos\left(\dfrac{I}{I_0} \right) = \arccos0.9 \approx 25\degree"

Answer. 1) 5, 2) 25 degrees.